A car mass 1400kg moves with a steady speed of 20 m/s on a straight rough road with its engine working at a constant rate of 36KW.
(a) Calculate the resistance R to the motion of the car if the road is horizontal.
Given that R is proportional to the square of the speed of the car,
(b) Show that R = 9v^2 /2 .
1
Expert's answer
2021-02-07T19:18:30-0500

"\text{2 forces act on the car along the X axis}"

"R - \text{resistance to movement}"

"F_m-\text{car motor power}"

"\text{the direction of these forces is opposite}"

"\text{the car moves at a constant speed according to Newton's first law,}"

"\text{the resultant of forces is 0}"

"F= F_m-R;F=0;R=F_m"

"a)P= F_m*v"

"F_m=\frac{P}{v}=\frac{36000}{20}=1800"

"R=F_m = 1800 N"

"b)R =k*v^2"

"k= \frac{R}{v^2}= \frac{1800}{20^2}= \frac{9}{2}"

"R= \frac{9v^2}{2}"

Answer:"a)R = 1800N" "b)R= \frac{9v^2}{2}"